I know the answer is no solution but i need to know how to work this problem please help
I+know+the+answer+is+no+solution+but+i+need+to+know+how+to+work+this+problem+please+help%0D%0A%0D%0A6+log(2,(2x-4))=7+log(2,(3x))

Answer by stanbon(48569)

6+log(2,(2x-4))=7+log(2,(3x))
Keep in mind, the logs are base two.
---------
log(2x-4)=1+log(3x)
log(2x-4)-log(3x)=1
log[(2x-4)/3x]=1
Since the logs are base 2 you get:
(2x-4)/3x=2^1
2x-4=6x
4x=-4
x=-1
Checking this in the original equation you get log(-3) which doesn't exist.
So, no solution.
Cheers,
Stan H.